OpenStudy (anonymous):

write -7y=8x-3 in slope intercept form. i got -8/7x-3/7=y. Did i do it right? Please help.

OpenStudy (anonymous):

Also -5/4x=25=y is what i got for 5x+4y=100

OpenStudy (anonymous):

*my bad -5/4x+25=y

OpenStudy (anonymous):

First one should be y=-8/7x+3/7